# UAV-Enabled ISAC Systems with Fluid Antennas

### v1

@@ III-C Target Sensing Model
where $\mathbf{X}_{t}[n]=\big[\mathbf{x}_{t}[1,n],\mathbf{x}_{t}[2,n],\ldots,\mathbf{x}_{t}[\tilde{L},n]\big]\in\mathbb{C}^{N_{t}\times\tilde{L}}$, and $\mathbf{N}_{r}[n]\in\mathbb{C}^{N_{r}\times\tilde{L}}$ denotes an AWGN matrix with zero mean and the variance of each its entry is $\sigma_{r}^{2}$. Assuming that $\tilde{L}$ to be sufficiently large, the covariance matrix of the transmitted signal can be approximated as
@@ III-E Three-Timescale Optimization Framework
Although frequent updates of the FA positions in each time slot could potentially enhance both communication and sensing performance, they also give rise to several practical limitations. First, overly frequent updates may incur response delays due to mechanical and control latency. Second, the continuous adjustment of FA positions increases the UAV’s energy consumption and control overhead, thereby reducing operational efficiency. Third, rapid position variations may induce non-ideal effects, such as Doppler frequency shifts and phase jitter, which lead to higher channel estimation errors and degraded communication quality.
To overcome these challenges, we propose a three-timescale optimization framework, in which the positions of the FA arrays are updated at a medium timescale, i.e., once in every $\mu$ time slots, rather than in every slot. As depicted in Fig. 2, the proposed framework operates over three timescales:
  • Long timescale (entire period): The UAV trajectory is globally optimized over the full entire period to account for mission objectives.
  • Medium timescale (time intervals): The FA array positions are updated once in each time interval to balance performance gains with energy consumption.
  • Short timescale (time slots): Beamforming is dynamically adjusted in every time slot to adapt to fast-varying channel conditions.
Here, $\tilde{\mathcal{I}}=\{1,\dots,\tilde{I}\}$ denotes the set of intervals, and $N=\mu\tilde{I}$ with $\mu\in\mathbb{Z}$.
Figure 2: Illustration of three-timescale framework.
@@ III-F Problem Formulation
In this work, we aim to develop a joint optimization framework that balances the trade-off between two competing objectives: minimizing the CRB on the angle estimation (or equivalently, maximizing the inverse of the CRB, as defined in (16)) and maximizing the ergodic sum rate of all users. To achieve this, we introduce weight factors $\xi_{c}$ and $\xi_{s}$, representing the relative priorities of communication performance and sensing accuracy, respectively, where $\xi_{c}+\xi_{s}=1$ and $\xi_{c},\xi_{s}\in[0,1]$. Accordingly, the objective function of the joint optimization problem is given by
@@
The objective is achieved by jointly optimizing the following variables: the UAV trajectory $\{\mathbf{q}[n]\}$, the beamforming design $\{\mathbf{w}_{m}[n],\mathbf{R}_{0}[n]\}$, the position of the transmit FA array $\{\mathbf{x}[n]\}$, and the position of the receive FA array $\{\mathbf{y}[n]\}$. Therefore, the optimization problem can be formulated as
@@
where constraints (18b) and (18c) describe the kinematic limitations of the UAV. Specifically, the UAV must depart from and return to predetermined initial and final locations, denoted by $\mathbf{q}_{\rm{I}}$ and $\mathbf{q}_{\rm{F}}$, respectively. Additionally, its velocity at each time slot is constrained by the maximum allowable value $V_{\rm{max}}$. Constraint (18d) indicates that the transmit power of the UAV must not exceed a predefined limit $P_{\max}$. Constraints (18e) and (18f) ensure that the positions of the FA arrays remain constant over the medium timescale. Linear inequality constraints (18g) and (18h) indicate that the distance between any two transmit (or receive) FAs must be greater than the minimum distance $d_{\rm{min}}$ to avoid antenna coupling, and all antennas are required to move within a one-dimensional segment $0,D_{\rm{FA}}]$. Specifically, the matrix $\mathbf{P}\in\mathbb{R}^{(N_{t}+1)\times N_{t}}$ and vector $\mathbf{l}_{\rm{p}}\in\mathbb{R}^{(N_{t}+1)\times 1}$ that define the constraints on the position of transmit FA array are given by
@@
The matrices $\mathbf{Q}$ and $\mathbf{l}_{\rm{q}}$ that define the corresponding constraints for the position of the receive FA array have a similar structure and are omitted for brevity.
@@ IV Proposed Solution
Due to the non-concavity of the objective function and the strong coupling among variables, problem (P1) is challenging to solve for a globally optimal solution. Furthermore, the UAV trajectory and the antenna positions are intricately embedded within the channel vector model. To address problem (P1), we first equivalently reformulate the problem by decomposing it into four subproblems. Subsequently, we employ an AO-based algorithm, iteratively optimizing the UAV trajectory, the beamforming design, the position of transmit FA array, and the position of the receive FA array, in sequence, to obtain a sub-optimal solution.
@@ IV-A Transmit Beamforming Optimization
Given the UAV trajectory $\{\mathbf{q}[n]\}$ and the positions of the FA arrays $\{\mathbf{x}[n],\mathbf{y}[n]\}$, the transmit beamforming optimization problem is formulated as
@@
$\small\tilde{R}_{m}[n]=\log_{2}\left(1+\frac{\zeta^{\rm{los}}_{m}[n]\left|\mathbf{\overline{h}}_{m}^{\mathsf{H}}[n]\mathbf{w}_{m}[n]\right|^{2}+\zeta^{\rm{nlos}}_{m}[n]\|\mathbf{w}_{m}[n]\|^{2}}{\sum\limits_{i=1,i\neq m}^{\mathcal{M}}\left(\zeta^{\rm{los}}_{m}[n]\left|\mathbf{\overline{h}}_{m}^{\mathsf{H}}[n]\mathbf{w}_{i}[n]\right|^{2}+\zeta^{\rm{nlos}}_{m}[n]\|\mathbf{w}_{i}[n]\|^{2}\right)+\zeta^{\rm{los}}_{m}[n]\mathbf{\overline{h}}_{m}^{\mathsf{H}}[n]\mathbf{R}_{0}[n]\mathbf{\overline{h}}_{m}[n]+\zeta^{\rm{nlos}}_{m}[n]\mathsf{tr}\left(\mathbf{R}_{0}[n]\right)+\sigma^{2}_{m}}\right).$ (22)
@@
$+\sum_{i=1,i\neq m}^{M}\left(\zeta^{\rm{los}}_{m}[n]\overline{\mathbf{h}}_{m}^{\mathsf{H}}[n]\mathbf{W}_{i}[n]\overline{\mathbf{h}}_{m}[n]+\zeta^{\rm{nlos}}_{m}[n]\mathsf{tr}\left(\mathbf{W}_{i}[n]\right)\right),$ (27)
@@
Specifically, $\tilde{R}^{\rm{III}}_{m}[n]$ in $\tilde{\mathcal{F}}_{1}$ is given by
@@
IV-B Position Optimization for Transmit FA Array
Based on Proposition 1, given the UAV trajectory $\{\mathbf{q}[n]\}$, the transmit beamforming $\{\mathbf{w}_{m}[n],\mathbf{R}_{0}[n]\}$ and the position of the receive FA array $\{\mathbf{y}[n]\}$, the position of the transmit FA array optimization problem is formulated as
@@ IV-B Position Optimization for Transmit FA Array
where
$\Psi_{m,p,q}[n]=\vartheta_{m}[n](x_{p}[n]-x_{q}[n])-\angle\left([\mathbf{W}_{m}[n]]_{p,q}\right).$ (37)
Since $\overline{\mathsf{h}}_{m}^{\mathrm{H}}[n]\mathbf{W}_{m}[n]\overline{\mathbf{h}}_{m}[n]$ is non-concave with respect to $\mathbf{x}[n]$, we construct a lower bound surrogate function to locally approximate it using its second-order Taylor expansion [38]. Specifically, for a given $a\in\mathbb{R}$, we can obtain
@@
$=\frac{1}{2}\mathbf{x}^{\mathsf{T}}[n]\mathbf{S}^{\rm{I}}_{m}[n]\mathbf{x}[n]+\left(\mathbf{t}_{m}^{\rm{I}}[n]\right)^{\mathsf{T}}\mathbf{x}[n]+u_{m}[n],$ (39)
@@
$\mathbf{S}^{\rm{I}}_{m}[n]=-2\vartheta_{m}^{2}[n]\left(\mathsf{diag}\left(\mathbf{V}_{m}^{\rm{I}}[n]\right)-\mathbf{V}_{m}^{\rm{II}}[n]\right),$ (40)
$\mathbf{V}_{m}^{\rm{I}}[n]=\sum_{q=1}^{N_{t}}\big[|[\mathbf{W}_{m}[n]]_{1,q}|,|[\mathbf{W}_{m}[n]]_{2,q}|,\ldots,|[\mathbf{W}_{m}[n]]_{N_{t},q}|\big]^{\mathsf{T}},\ \mathbf{V}_{m}^{\rm{II}}[n]=\left(|[\mathbf{W}_{m}[n]]_{p,q}|\right)_{p,q},$ (41)
@@
Since $\mathbf{S}^{\rm{I}}_{m}[n]$ is negative semi-definite, constraint (44) is convex in $\mathbf{x}[n]$. The complete proof is provided in Appendix D.
@@
Since $\mathbf{S}^{\rm{I}}_{m}[n]$ is positive semi-definite, constraint (50) is concave in $\mathbf{x}[n]$.
Similarly, we can construct lower bound and upper bound surrogate functions for $\mathbf{a}^{\mathsf{H}}[n]\mathbf{R}_{\mathbf{x}}[n]\mathbf{a}[n]$ in $\frac{1}{\text{CRB}(\theta_{\rm{T}}[n])}$, as well as for $\overline{\mathbf{h}}_{m}^{\mathsf{H}}[n]\mathbf{W}_{i}[n]\overline{\mathbf{h}}_{m}[n]$ and $\overline{\mathbf{h}}_{m}^{\mathsf{H}}[n]\mathbf{R}_{0}[n]\overline{\mathbf{h}}_{m}[n]$ in $F_{m}[n]$. Accordingly, we derive the associated constraints based on these surrogate functions, which take the following general forms,
@@
where the index $j={{\rm{II}}(i)},{\rm{III}},{\rm{IV}}$ corresponds to the terms $\overline{\mathbf{h}}_{m}^{\mathsf{H}}[n]\mathbf{W}_{i}[n]\overline{\mathbf{h}}_{m}[n]$, $\overline{\mathbf{h}}_{m}^{\mathsf{H}}[n]\mathbf{R}_{0}[n]\overline{\mathbf{h}}_{m}[n]$, and $\mathbf{a}^{\mathsf{H}}[n]\mathbf{R}_{\mathbf{x}}[n]\mathbf{a}[n]$, respectively.
@@
IV-C Position Optimization for Receive FA Array
Obviously, the ergodic sum rate of all users is independent of the position of the receive FA array. Therefore, given the UAV trajectory $\{\mathbf{q}[n]\}$, the transmit beamforming $\{\mathbf{w}_{m}[n],\mathbf{R}_{0}[n]\}$, and the position of the transmit FA array $\{\mathbf{x}[n]\}$, the position of the receive FA array optimization problem can be formulated as
@@ IV-C Position Optimization for Receive FA Array
In statistical theory, given a set of observed data $\{y_{1},y_{2},\cdots,y_{n}\}$, the sample mean and the total sum of squares (TSS) are defined as $\bar{y}=\frac{1}{n}\sum_{i=1}^{n}y_{i}$ and $\text{TSS}=\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$, respectively. Based on this definition, the quadratic form $\mathbf{y}[n]^{\mathsf{T}}\left(\mathbf{I}_{N_{r}}-\frac{1}{N_{r}}\mathbf{1}_{N_{r}}\mathbf{1}_{N_{r}}^{\mathsf{T}}\right)\mathbf{y}[n]$ is equivalent to the TSS of the sequence $\mathbf{y}[n]$, and is equal to the sample variance scaled by a factor of $(N_{r}-1)$. Therefore, the optimization objective corresponds to the weighted sum of the TSS across all time slots.
@@ IV-D UAV Trajectory Optimization
Given the transmit beamforming $\{\mathbf{w}_{m}[n],\mathbf{R}_{0}[n]\}$ and the positions of the FA arrays $\{\mathbf{x}[n],\mathbf{y}[n]\}$, the UAV trajectory optimization problem is formulated as
@@
Next, we address the non-concavity of the new optimization objective $\mathcal{F}_{1}^{(l)}$. For (66), it can be rewritten as

### v2

@@ III-C Target Sensing Model
where $\mathbf{X}_{t}[n]=\big[\mathbf{x}_{t}[1,n],\mathbf{x}_{t}[2,n],\ldots,\mathbf{x}_{t}[\tilde{L},n]\big]\in\mathbb{C}^{N_{t}\times\tilde{L}}$, and $\mathbf{N}_{r}[n]\in\mathbb{C}^{N_{r}\times\tilde{L}}$ denotes an AWGN matrix with zero mean and the variance of each its entry is $\sigma_{r}^{2}$. Assuming that $\tilde{L}$ is sufficiently large, the covariance matrix of the transmitted signal can be approximated as
@@ III-E Three-Timescale Optimization Framework
To balance performance gains and practical implementation constraints, we adopt a three-timescale optimization framework. Although frequently updating the FA positions potentially enhance system performance, excessively frequent adjustments may incur additional control latency, increase energy consumption, and introduce non-ideal effects such as Doppler shifts and phase instability.
To address these issues, the system design is divided into three hierarchical timescales, as illustrated in Fig. 2. Specifically, the UAV trajectory is optimized at a long timescale over the entire mission period to account for large-scale propagation characteristics. The FA positions are updated at a medium timescale, i.e., in every $\mu$ time slots, to balance spatial flexibility and mechanical overhead. Meanwhile, the transmit beamforming is optimized at a short timescale, i.e., in every time slot, to adapt to instantaneous channel variations. In Fig. 2, $\tilde{\mathcal{I}}=\{1,\dots,\tilde{I}\}$ denotes the set of intervals, and $N=\mu\tilde{I}$ with $\mu\in\mathbb{Z}$.
Figure 2: Illustration of three-timescale optimization framework.
@@ III-F Problem Formulation
In this work, we aim to develop a joint optimization framework that balances the trade-off between two competing objectives: minimizing the CRB on the angle estimation (or equivalently, maximizing the inverse of the CRB, as defined in (16)) and maximizing the ergodic sum rate of all users. To achieve this, we introduce weight factors $\xi_{c}$ and $\xi_{s}$, representing the relative priorities of communication performance and sensing accuracy, respectively. Accordingly, the objective function of the joint optimization problem is given by
@@
The objective is achieved by jointly optimizing the following variables: the UAV trajectory $\{\mathbf{q}[n]\}$, the beamforming design $\{\mathbf{w}_{m}[n],\mathbf{R}_{0}[n]\}$, the transmit FA positions $\{\mathbf{x}[n]\}$, and the receive FA positions $\{\mathbf{y}[n]\}$. Therefore, the optimization problem can be formulated as
@@
where constraints (18b) and (18c) describe the kinematic limitations of the UAV. Specifically, the UAV must depart from and return to the predetermined initial and final locations, denoted by $\mathbf{q}_{\rm{I}}$ and $\mathbf{q}_{\rm{F}}$, respectively. Additionally, its velocity in each time slot is constrained by the maximum allowable value $V_{\rm{max}}$. Constraint (18d) indicates that the transmit power of the UAV must not exceed a predefined limit $P_{\max}$. Constraints (18e) and (18f) ensure that the FA positions remain constant over the medium timescale. Constraints (18g) and (18h) indicate that the distance between any two transmit/receive FAs must be greater than the minimum distance $d_{\rm{min}}$ to avoid antenna coupling, and all antennas are required to move within a one-dimensional segment $0,D_{\rm{FA}}]$. Specifically, the matrix $\mathbf{P}\in\mathbb{R}^{(N_{t}+1)\times N_{t}}$ and vector $\mathbf{l}_{\rm{p}}\in\mathbb{R}^{(N_{t}+1)\times 1}$ that define the constraints on the transmit FA positions are given by
@@
The matrices $\mathbf{Q}$ and $\mathbf{l}_{\rm{q}}$ that define the corresponding constraints for the receive FA positions have a similar structure and are omitted for brevity.
@@ IV Proposed Solution
Due to the non-concavity of the objective function and the strong coupling among variables, problem (P1) is challenging to solve for a globally optimal solution. Furthermore, the UAV trajectory and the antenna positions are intricately embedded within the channel vector model. To address these, we first equivalently reformulate the problem by decomposing it into four subproblems. Subsequently, we employ an AO-based algorithm, iteratively optimizing the UAV trajectory, the beamforming design, the transmit FA positions, and the receive FA positions, in sequence, to obtain a sub-optimal solution.
@@ IV-A Transmit Beamforming Optimization
Given the UAV trajectory $\{\mathbf{q}[n]\}$ and the FA positions $\{\mathbf{x}[n],\mathbf{y}[n]\}$, the optimization problem for the transmit beamforming is formulated as
@@
$\small\tilde{R}_{m}[n]=\log_{2}\left(1+\frac{\zeta^{\rm{los}}_{m}[n]\left|\mathbf{\overline{h}}_{m}^{\mathsf{H}}[n]\mathbf{w}_{m}[n]\right|^{2}+\zeta^{\rm{nlos}}_{m}[n]\|\mathbf{w}_{m}[n]\|^{2}}{\sum\limits_{i\neq m}^{\mathcal{M}}\left(\zeta^{\rm{los}}_{m}[n]\left|\mathbf{\overline{h}}_{m}^{\mathsf{H}}[n]\mathbf{w}_{i}[n]\right|^{2}+\zeta^{\rm{nlos}}_{m}[n]\|\mathbf{w}_{i}[n]\|^{2}\right)+\zeta^{\rm{los}}_{m}[n]\mathbf{\overline{h}}_{m}^{\mathsf{H}}[n]\mathbf{R}_{0}[n]\mathbf{\overline{h}}_{m}[n]+\zeta^{\rm{nlos}}_{m}[n]\mathsf{tr}\left(\mathbf{R}_{0}[n]\right)+\sigma^{2}_{m}}\right).$ (22)
@@
$+\sum_{i\neq m}^{M}\left(\zeta^{\rm{los}}_{m}[n]\overline{\mathbf{h}}_{m}^{\mathsf{H}}[n]\mathbf{W}_{i}[n]\overline{\mathbf{h}}_{m}[n]+\zeta^{\rm{nlos}}_{m}[n]\mathsf{tr}\left(\mathbf{W}_{i}[n]\right)\right),$ (27)
@@
Specifically, $\tilde{R}^{\rm{III}}_{m}[n]$ is given by
@@
IV-B Transmit FA Position Optimization
Based on Proposition 1, given the UAV trajectory $\{\mathbf{q}[n]\}$, the transmit beamforming $\{\mathbf{w}_{m}[n],\mathbf{R}_{0}[n]\}$, and the receive FA positions $\{\mathbf{y}[n]\}$, the optimization problem for the transmit FA positions is formulated as
@@ IV-B Transmit FA Position Optimization
where $\Psi_{m,p,q}[n]=\vartheta_{m}[n](x_{p}[n]-x_{q}[n])-\angle\left([\mathbf{W}_{m}[n]]_{p,q}\right).$ Since $\overline{\mathsf{h}}_{m}^{\mathrm{H}}[n]\mathbf{W}_{m}[n]\overline{\mathbf{h}}_{m}[n]$ is non-concave with respect to $\mathbf{x}[n]$, we construct a lower bound surrogate function to locally approximate it using its second-order Taylor expansion [38]. Specifically, for a given $a\in\mathbb{R}$, we can obtain
@@
$=\frac{1}{2}\mathbf{x}^{\mathsf{T}}[n]\mathbf{S}^{\rm{I}}_{m}[n]\mathbf{x}[n]+\left(\mathbf{t}_{m}^{\rm{I}}[n]\right)^{\mathsf{T}}\mathbf{x}[n]+u_{m}^{\rm{I}}[n],$ (38)
@@
$\mathbf{S}^{\rm{I}}_{m}[n]=-2\vartheta_{m}^{2}[n]\left(\mathsf{diag}\left(\mathbf{V}_{m}^{\rm{I}}[n]\right)-\mathbf{V}_{m}^{\rm{II}}[n]\right),\mathbf{V}_{m}^{\rm{I}}[n]=\sum_{q=1}^{N_{t}}\big[|[\mathbf{W}_{m}[n]]_{1,q}|,\ldots,|[\mathbf{W}_{m}[n]]_{N_{t},q}|\big]^{\mathsf{T}},\ \mathbf{V}_{m}^{\rm{II}}[n]=\left(|[\mathbf{W}_{m}[n]]_{p,q}|\right)_{p,q},$ (39)
@@
Since $\mathbf{S}^{\rm{I}}_{m}[n]$ is negative semi-definite, constraint (42) is convex with respect to $\mathbf{x}[n]$. The complete proof is provided in Appendix D.
@@
Since $\mathbf{S}^{\rm{I}}_{m}[n]$ is positive semi-definite, constraint (48) is concave with respect to $\mathbf{x}[n]$.
Similarly, we can construct the lower bound and upper bound surrogate functions for $\mathbf{a}^{\mathsf{H}}[n]\mathbf{R}_{\mathbf{x}}[n]\mathbf{a}[n]$ in $\frac{1}{\text{CRB}(\theta_{\rm{T}}[n])}$, as well as for $\overline{\mathbf{h}}_{m}^{\mathsf{H}}[n]\mathbf{W}_{i}[n]\overline{\mathbf{h}}_{m}[n]$ and $\overline{\mathbf{h}}_{m}^{\mathsf{H}}[n]\mathbf{R}_{0}[n]\overline{\mathbf{h}}_{m}[n]$ in $F_{m}[n]$. Accordingly, we derive the associated constraints based on these surrogate functions, which take the following general forms,
@@
where the index $j={{\rm{II}}(i)},{\rm{III}},$ and $\rm{IV}$} corresponds to the terms $\overline{\mathbf{h}}_{m}^{\mathsf{H}}[n]\mathbf{W}_{i}[n]\overline{\mathbf{h}}_{m}[n]$, $\overline{\mathbf{h}}_{m}^{\mathsf{H}}[n]\mathbf{R}_{0}[n]\overline{\mathbf{h}}_{m}[n]$, and $\mathbf{a}^{\mathsf{H}}[n]\mathbf{R}_{\mathbf{x}}[n]\mathbf{a}[n]$, respectively.
@@
IV-C Receive FA Position Optimization
Obviously, the ergodic sum rate of all users is independent of the receive FA positions. Therefore, given the UAV trajectory $\{\mathbf{q}[n]\}$, the transmit beamforming $\{\mathbf{w}_{m}[n],\mathbf{R}_{0}[n]\}$, and the transmit FA positions $\{\mathbf{x}[n]\}$, the optimization problem for the receive FA positions can be formulated as
@@ IV-C Receive FA Position Optimization
In statistical theory, given a set of observed data $\{y_{1},y_{2},\cdots,y_{n}\}$, the sample mean and the total sum of squares (TSS) are defined as $\bar{y}=\frac{1}{n}\sum_{i=1}^{n}y_{i}$ and $\text{TSS}=\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$, respectively. Based on these definitions, the quadratic form $\mathbf{y}[n]^{\mathsf{T}}\left(\mathbf{I}_{N_{r}}-\frac{1}{N_{r}}\mathbf{1}_{N_{r}}\mathbf{1}_{N_{r}}^{\mathsf{T}}\right)\mathbf{y}[n]$ is equivalent to the TSS of the sequence $\mathbf{y}[n]$, and is equal to the sample variance scaled by a factor of $(N_{r}-1)$. Therefore, the objective function corresponds to the weighted sum of the TSS across all time slots.
@@ IV-D UAV Trajectory Optimization
Given the transmit beamforming $\{\mathbf{w}_{m}[n],\mathbf{R}_{0}[n]\}$ and the FA positions $\{\mathbf{x}[n],\mathbf{y}[n]\}$, the optimization problem for the UAV trajectory is formulated as
@@
Next, we address the non-concavity of the new objective function $\mathcal{F}_{1}^{(l)}$. For (64), it can be rewritten as